OpenStudy (anonymous):

Evaluate the following limit: \[\lim_{x \rightarrow 0-} \frac{e^{1/x}}{x}\]

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@amistre64 @AccessDenied @JamesJ @SmoothMath (this is a different question :p)

OpenStudy (anonymous):

I know this is easy, but I do not know how to proceed with this

OpenStudy (anonymous):

I believe L'hopital's rule applies... but I'm not sure if it's usable for one sided limits.

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

Well then cool. Do you see that the top and bottom both approach 0?

OpenStudy (anonymous):

From the left, that is. From the right, the top would approach infinity.

OpenStudy (anonymous):

How will it work? You get: \[\lim_{x \rightarrow 0-} \frac{e^{1/x}*-1/x^2}{1}\]

OpenStudy (anonymous):

Which will then go to the denominator, and then you do it and over and over and over

OpenStudy (anonymous):

I think there is no need for L'H Rule here

OpenStudy (anonymous):

I think you're right.

OpenStudy (anonymous):

But then wait, if I sub in close to 0, then x becomes 0, but then e^{1/x} goes to infinity (v.large number) that goes into the denominator

OpenStudy (anonymous):

so we get: \[\frac{1}{\infty*0}\]

OpenStudy (anonymous):

You have a very large number multiplied by 0. That is 0 then?

OpenStudy (anonymous):

\[\infty * 0\]

OpenStudy (anonymous):

b/c: \[\frac{1}{\infty}*\frac{1}{0}\]

OpenStudy (anonymous):

No no. It's still as x approaches 0 from the left, which means e^(1/x) approaches 0.

OpenStudy (unklerhaukus):

i dont follow the argument

OpenStudy (experimentx):

must be zero .... exponential function always win over polynomial function ...

OpenStudy (zarkon):

let \(u=-\frac{1}{x}\) then \[\lim_{x \to 0^{-}} \frac{e^{1/x}}{x}=\lim_{u\to\infty}-\frac{e^{-u}}{1/u}\] \[=-\lim_{u\to\infty}\frac{u}{e^{u}}\] now apply L'Hospital's rule

OpenStudy (anonymous):

Seems legit.

OpenStudy (unklerhaukus):

the answer is \(0^-\)right?

OpenStudy (unklerhaukus):

or just \(0\)

OpenStudy (zarkon):

the answer is zero

OpenStudy (anonymous):

And to test that, you can check values increasingly closer to 0 from the left. Now, this isn't fool-proof. Some functions will mislead you if you try this kind of test. It usually works though.